OpenStudy (anonymous):

How many units apart are the points (8, 11) and (-1, -1)?

OpenStudy (anonymous):

@DannyRomero19

OpenStudy (anonymous):

Use distance formula: \[Distance = \sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2}\]

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

\[x_2 = -1, x_1 = 8, y_2 = -1, y_1 = 11\]

OpenStudy (anonymous):

Try to plug in the values and find the distance..

OpenStudy (anonymous):

so the distance is = 11 units totALL?

OpenStudy (anonymous):

How ??

OpenStudy (anonymous):

idk how to do it can you just tell me the total units?

OpenStudy (anonymous):

Your work is to plug in the values and then find the distance.. Try it you can do..

OpenStudy (anonymous):

16

OpenStudy (anonymous):

You got \(\sqrt{256}\) ??

OpenStudy (anonymous):

Take the square root properly.. You are close to it..

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

15.8

OpenStudy (anonymous):

how can you show me what you did?? So that I can tell you where you are wrong..

OpenStudy (anonymous):

I think you are given with answer choices.. Right ??

OpenStudy (anonymous):

i was not given choices :(

OpenStudy (anonymous):

\[Distance = \sqrt{(-1 -8)^2 + (-1 - 11)^2} \implies \sqrt{9^2 + 12^2} \implies \sqrt{225}\] Now can you find its value??

OpenStudy (anonymous):

15

OpenStudy (anonymous):

Yep..